# Constrained Style Learning from Imperfect Demonstrations under Task Optimality

### v1

@@ Abstract
Learning from demonstration has proven effective in robotics for acquiring natural behaviors, such as stylistic motions and lifelike agility, particularly when explicitly defining style-oriented reward functions is challenging. Synthesizing stylistic motions for real-world tasks usually requires balancing task performance and imitation quality. Existing methods generally depend on expert demonstrations closely aligned with task objectives. However, practical demonstrations are often incomplete or unrealistic, causing current methods to boost style at the expense of task performance. To address this issue, we propose formulating the problem as a constrained Markov Decision Process (CMDP). Specifically, we optimize a style-imitation objective with constraints to maintain near-optimal task performance. We introduce an adaptively adjustable Lagrangian multiplier to guide the agent to imitate demonstrations selectively, capturing stylistic nuances without compromising task performance. We validate our approach across multiple robotic platforms and tasks, demonstrating both robust task performance and high-fidelity style learning. On ANYmal-D hardware we show a drop in mechanical energy and a more agile gait pattern, showcasing real-world benefits.
@@ 2 Related Work
Despite their success, these approaches implicitly assume that demonstration data will always help or at least not hinder task performance. In practice, this assumption does not hold: high-performing policies often require generalization to diverse terrains and command ranges [24], while demonstrations may be limited, misaligned, or collected under drastically different conditions. Blindly following such data can negatively impact performance, especially when task requirements conflict with stylistic cues. Our work builds upon these insights and proposes a principled framework to reconcile this conflict. By adopting a CMDP-based policy optimization framework and using a self-adjustable Lagrangian multiplier to adjust imitation weight, we allow the agent to selectively benefit from imperfect demonstrations without compromising its ability to learn robust task policies.
@@ 5 Experiments and Results
As shown in Fig. 2, ConsMimic consistently achieves an effective balance between task performance and imitation quality. Specifically, in Franka-Reach, ConsMimic obtains a high task reward (0.78) and demonstrates significantly better imitation quality compared to the task-only baseline. In quadruped tasks, particularly Anymal-Lateral where style-task misalignment is prominent, ConsMimic achieves high task performance close to the task-only baseline while substantially surpassing fixed-weight baselines in imitation quality, highlighting its capacity to effectively leverage stylistic cues without compromising task objectives. For humanoid tasks, fixed-weight baselines either sacrifice task completion capability or experience a significant decline in stylistic imitation. For instance, demonstrates strong imitation but fails in complex scenarios like GR1-Stair and GR1-Stepstone, whereas maintains task performance but consistently scores lower in imitation quality. ConsMimic’s adaptive strategy consistently delivers better generalization by dynamically balancing style imitation and task requirements, enabling stable training and robust performance under realistic task-demo misalignments. We provide details of DTW calculation in Appendix C.
@@
Task | Ours (w/o sym aug) | Ours (w/ sym aug) |
GR1-Flat | ||
GR1-Stair | ||
@@
Q4. ConsMimic on Real-World Tasks. We further validate our framework on the ANYmal-D quadruped. The policy, trained on the Anymal-Forward task, is deployed on the ANYmal-D hardware in a zero-shot manner. The robot is commanded to move forward at a speed of 2 m/s and then return at the same speed. It completes 8 episodes in total while traversing an identical distance in each episode.
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.

### v2

@@ Abstract
Learning from demonstration has proven effective in robotics for acquiring natural behaviors, such as natural motions and lifelike agility, particularly when explicitly defining style-oriented reward functions is challenging. Synthesizing stylistic motions for real-world tasks usually requires balancing task performance and imitation quality. Existing methods generally depend on expert demonstrations closely aligned with task objectives. However, practical demonstrations are often incomplete or unrealistic, causing current methods to boost style at the expense of task performance. To address this issue, we propose formulating the problem as a constrained Markov Decision Process (CMDP). Specifically, we optimize a style-imitation objective with constraints to maintain near-optimal task performance. We introduce an adaptively adjustable Lagrangian multiplier to guide the agent to imitate demonstrations selectively, capturing stylistic nuances without compromising task performance. We validate our approach across multiple robotic platforms and tasks, demonstrating both robust task performance and high-fidelity style learning. On ANYmal-D hardware we show a drop in mechanical energy and a more agile gait pattern, showcasing real-world benefits.
@@ 2 Related Work
Despite their success, these approaches implicitly assume that demonstration data will always help or at least not hinder task performance. In practice, this assumption does not hold: high-performing policies often require generalization to diverse terrains and command ranges [24], while demonstrations may be limited, misaligned, or collected under drastically different conditions. Blindly following such data can negatively impact performance, especially when task requirements conflict with stylistic cues. Our work builds upon these insights and proposes a principled framework to reconcile this conflict. By adopting a CMDP-based policy optimization framework and using a self-adjustable Lagrangian multiplier to adjust imitation weight, we allow the agent to selectively learn from imperfect demonstrations without compromising its ability to learn robust task policies.
@@ 5 Experiments and Results
As shown in Fig. 2, ConsMimic consistently achieves an effective balance between task performance and imitation quality. Specifically, in Franka-Reach, ConsMimic obtains a high task reward (0.78) and demonstrates significantly better imitation quality compared to the task-only baseline. In quadruped tasks, particularly Anymal-Lateral where style-task misalignment is prominent, ConsMimic achieves high task performance close to the task-only baseline while substantially surpassing fixed-weight baselines in imitation quality, highlighting its capacity to effectively leverage stylistic cues without compromising task objectives. For humanoid tasks, fixed-weight baselines either sacrifice task completion capability or experience a significant decline in stylistic imitation. For instance, demonstrates strong imitation but fails in complex scenarios like GR1-Stair and GR1-Stepstone, whereas maintains task performance but consistently scores lower in imitation quality. ConsMimic’s adaptive strategy consistently delivers better generalization by dynamically balancing style imitation and task requirements, enabling stable training and robust performance under realistic task-demo misalignments. We provide details for DTW calculation in Appendix C.
@@
| Task |
| GR1-Flat | ||||||
| GR1-Stair | ||||||
@@
Q4. ConsMimic on Real-World Tasks. We further validate our framework on the ANYmal-D quadruped. The policy, trained on the Anymal-Forward task, is deployed on the ANYmal-D hardware in a zero-shot manner. The robot is commanded to move forward at a speed of 2 m/s and then return at the same speed. It completes 8 rounds with identical distances.
@@ Acknowledgments
This research was supported by the ETH AI Center and the Swiss National Science Foundation through the National Centre of Competence in Automation (NCCR automation).